OpenStudy (anonymous):

Solve for c. 2abc + d = 3

OpenStudy (anonymous):

We want to isolate c on one side. We could start by subtracting d from both sides to get: \[2abc=3-d\] Finally, divide both sides by 2ab: \[c=(3-d)/(2ab)\]

OpenStudy (anonymous):

2abc = 3-d \[c = \left(\begin{matrix}3-d \\ 2ab\end{matrix}\right)\]

OpenStudy (anonymous):

Makes sense!! Thank you!!